OpenStudy (anonymous):

Okay younglings and fetuses.. Answer this question and win a medal , a fan , and Schrute buck.. y = log x If y = 10, then what is x?

OpenStudy (anonymous):

NAY! ^

OpenStudy (anonymous):

\[\eqalign{ &y=\log_ax \\ &10=log_ax \\ &a^{10}=a^{log_ax} \\ &a^{10}=x }\]

OpenStudy (anonymous):

\(a\) is the base of the logarithm...since you are just mentioning \(\log x\) I assume the base is 10 so we have: \[x=10^{10}\]

OpenStudy (anonymous):

Here is your Schrute buck.

OpenStudy (anonymous):

Well gee, thanks haha

OpenStudy (anonymous):

In general for having: \[y=log_ax\] Then: \[x=a^y\]

OpenStudy (anonymous):

Mhm well most of you're 1st answers were correct. I got y = log x 10 = log x 10 = 10^x x = 1

zepdrix (zepdrix):

Ooo nice one keith! :O Now you only need 999 more schrute bucks for an extra 5 minute lunch break!

OpenStudy (anonymous):

Exactly ! ^